# I Introduction

### v1

@@ V-D1 Scene Generation
To assess the diversity of scenario generation, we conducted three subsets of tests on MinADE and MinFDE for different agent types. The results, presented in the Table I, demonstrate that our approach achieves good predictive performance for various agents and can approximate the ground truth. In contrast, plain TrafficGen performs well for vehicle agents but exhibits significant performance degradation for pedestrians and bicycles due to its reliance on vehicle data, which oversimplifies the problem and reduces model generalizability. To ensure fairness, we trained trafficgen on a mixed dataset, which also yielded promising results on some metrics. This highlights the importance of establishing a framework for different agent types. DragTraffic outperforms other models on both vehicle and pedestrian datasets, thanks to its MOE structure, which enhances its generalizability. However, we observe that DragTraffic does not perform as well on the Cyclist dataset, possibly because the proportion of cyclists in the scene data is too small. In Table II, we present the results of the Interaction Reasoning Evaluation with sampling intervals of 3s, 6s, and 9s. We collected 300 scenes with high interactivity between different agents for evaluation, including 200 scenes related to pedestrians and 100 scenes related to cyclists. The performance of Dragtraffic under the three sampling conditions is superior to the baseline. Interestingly, we observe that the SCR varies with the sampling interval, which is contrary to the results obtained in Trafficgen. However, we believe this is reasonable because multiple rollouts can introduce more cumulative errors.
@@ V-D2 Scene Editing & Inpaiting
Figure 3 demonstrates the quality of scenes generated by DragTraffic based on existing data. We use the current frame information in the dataset and the information of the next nine seconds as conditions. The left-turn and right-turn scenes at the intersection where the most interactions occur are respectively shown in columns (a) and (b) of figure 3. These scenes reflect courtesy and competition for right of way among different agents. For example, in (a), Agent No. 9 and Agent No. 1 engage in a fierce competition for the right of way, while Agent No. 0 gives way. This shows that DragTraffic can simulate traffic participants in different situations under highly dynamic and complex traffic intersections to make reasonable, smooth, and realistic future actions, reflecting real-world characteristics. Next, we verify that this capability is not limited to simple log replay in more complex scenarios (c) and (d). We focus on agent No. 1, which is waiting at an intersection for the motorcycle in front to start. However, the movement of a large number of pedestrians around interferes with the decision-making of agent No. 1, which has a conservative driving style. As a result, it continues to watch the movements of pedestrians even when it is already driving to the intersection, leading to the phenomenon of robot freezing, which is common in the fields of autonomous driving and robotics. To address this issue, we design similar scenarios by setting the condition information for agent No. 1 (100 meters in front of it, longitudinal speed of 20m/s, and horizontal speed of -2m/s) to inspire more proactive behavior. As shown in (d), the results generated by DragTraffic are subject to our good scene editing control. Note that we only show the control of agent No. 1 for the convenience of explanation, but in fact, DragTraffic allows us to control the generation of multiple agents simultaneously. Unlike other methods of controlling generation (such as complex optimization constraints and loss design), these manipulation processes only require simple dragging or typing interactions, demonstrating the superiority of this framework in controlling degrees of freedom.
@@ VI Future Work
Although our proposed framework has the potential to generate realistic and diverse traffic scenarios, there are still areas that can be improved. For instance, we can use post-processing or sampling techniques to ensure that the generated driving actions meet dynamic constraints. Another promising direction for future work is to develop a more user-friendly interaction approach, such as incorporating large language models for text to motion generation. Additionally, due to space limitations, we were only able to present a portion of the editing potential. We encourage readers to explore more scene generation possibilities using Dragtraffic.
@@ VII CONCLUSIONS
In this paper, we presented DragTraffic, a generalized, point-based, and controllable traffic scene generation framework based on conditional diffusion. Our framework addresses the lack of controllability, accuracy, and versatility in existing scene generation methods, enabling non-experts to generate a variety of realistic driving scenarios for different types of traffic agents. We achieve this through an adaptive mixture expert architecture that uses a regression model to provide a general initial solution and a refinement process based on the conditional diffusion model to ensure diversity. The denoising process also introduces user-customized specific context through cross-attention, ensuring high controllability. Our experiments on a real-world driving dataset demonstrate that DragTraffic outperforms existing methods in terms of authenticity, diversity, and freedom. We believe that our framework can significantly contribute to the evaluation and training of autonomous driving systems by providing diverse and scalable corner cases.

### v2

@@ V-D1 Scene Generation
To assess the diversity of scenario generation, we conducted three subsets of tests on MinADE and MinFDE for different agent types. The results, presented in the Table I, demonstrate that our approach achieves good predictive performance for various agents and can approximate the ground truth. In contrast, plain TrafficGen performs well for vehicle agents but exhibits significant performance degradation for pedestrians and bicycles due to its reliance on vehicle data, which oversimplifies the problem and reduces model generalizability. To ensure fairness, we trained trafficGen on a mixed dataset, which also yielded promising results on some metrics. This highlights the importance of establishing a framework for different agent types. DragTraffic outperforms other models on both vehicle and pedestrian datasets, thanks to its MoE structure, which enhances its generalizability. However, we observe that DragTraffic does not perform well on the Cyclist dataset, possibly because the proportion of cyclists in the scene data is too small. In Table II, we present the results of the Interaction Reasoning Evaluation with sampling intervals of 3s, 6s, and 9s. We collected 300 scenes with high interactivity between different agents for evaluation, including 200 scenes related to pedestrians and 100 scenes related to cyclists. The performance of DragTraffic under the three sampling conditions is superior to the baseline. Interestingly, we observe that the SCR varies with the sampling interval, which is contrary to the results obtained in TrafficGen[8]. However, we believe this is reasonable because multiple rollouts can introduce more cumulative errors.
@@ V-D2 Scene Editing & Inpaiting
Fig. 3 demonstrates the quality of scenes generated by DragTraffic based on existing data. We use the information of current frame and the 90th frame in the dataset as conditions. The left-turn and right-turn scenes at the intersection where the most interactions occur are respectively shown in columns (a) and (b). These scenes reflect courtesy and competition for right of way among different agents. For example, in (a), Agent 9 and Agent 11 engage in a fierce competition for the right of way, while Agent 0 gives way. This shows that DragTraffic can simulate traffic participants in different situations under highly dynamic and complex traffic intersections to make reasonable, smooth, and realistic future actions, reflecting real-world characteristics. Next, we verify that this capability is not limited to simple log replay in more complex scenarios (c) and (d). We focus on agent 1, which is waiting at an intersection for the motorcycle in front to start. However, the movement of a large number of pedestrians around interferes with the decision-making of agent 1, which has a conservative driving style. As a result, it continues to watch the movements of pedestrians even when the motorcycle is already driving to the intersection, leading to the phenomenon of robot freezing, which is common in the fields of autonomous driving and robotics. To address this issue, we design similar scenarios by setting the condition information for agent 1 (100 meters in front of it, longitudinal speed of 20m/s, and latitudinal speed of -2m/s) to encourage more proactive behavior. As shown in (d), the results generated by DragTraffic demonstrate effective control over scene editing. Note that we only show the control of agent 1 for the convenience of explanation, but in fact, DragTraffic allows us to control the generation of multiple agents simultaneously. Unlike other generation control methods that rely on complex optimization constraints and loss designs, our generation processes require only simple dragging or typing interactions, showcasing the superiority of this framework.
@@ VI Future Work
While our proposed framework can generate realistic and diverse traffic scenarios, there are still areas for improvement. For example, we can implement post-processing or sampling techniques to ensure that the generated driving actions adhere to dynamic constraints. Another promising avenue for future work is to develop a multi-round generation process incorporating human feedback.
@@ VII CONCLUSIONS
In this paper, we introduced DragTraffic, a generalized, interactive, and controllable traffic scene generation framework based on conditional diffusion. Our framework addresses the limitations of existing scene generation methods in terms of controllability, accuracy, and versatility, enabling non-experts to create a variety of realistic driving scenarios for different types of traffic agents. We achieve this using an adaptive mixture expert architecture with a regression model for initial solutions, followed by refinement through the conditional diffusion model to ensure diversity. The denoising process incorporates user-customized context via cross-attention, enhancing controllability. The qualitative and quantitative results on a real-world driving dataset demonstrate the effectiveness of DragTraffic, which can significantly aid in the evaluation and training of autonomous driving systems by providing diverse and scalable corner cases.